250ml of "x" molarity sol. and 500ml of "y" molarity sol. of any solute are mixed and diluted to 2L to produce a final conc. of 1.6 Molarity. Calculate x and y if x/y= 5/4.

Grade:11

1 Answers

Manish kumar
4 years ago
According to dilution of molarity 
M1V1 +M2V2 +.......+MnVn =MrVr
250X +500Y=2000×1.6
=250X +500Y =200×16
=250X +5/4×500X   (Because 5/4=x/y given )
975x =3200
x =3200/975 =3.2
then, x/y =5/4.  Put the value of x and get 
Y=4
HOPE It is helpful for you 
975x